# Tagged particle fluctuations for TASEP with dynamics restricted by a moving wall

### v1

@@ 3.2 Asymptotic results
$\mathbb{P}(N(t\downarrow 0)>\kappa t^{1/3})=\mathbb{P}(|x_{N(t\downarrow 0)}(0)|\geq\kappa t^{1/3})\leq Ce^{-c\kappa},$ (3.34)
$\mathbb{P}(x_{P}^{\rho_{-}}(t)\leq(1-2\rho_{-})t-\kappa t^{2/3})\geq 1-Ce^{-c\kappa},$
@@
$Z_{\kappa,t}\coloneqq\#\{\textrm{particles in }x^{\rho_{-}}(0)\textrm{ at sites in }\{-\tfrac{1}{2}\kappa t^{2/3},\dots,-1\}\}\sim\textrm{Bin}(\tfrac{1}{2}\kappa t^{2/3},\rho_{-}).$ (3.36)
@@
with the same probability as $P(t\downarrow 0)>N(t\downarrow 0)$.
@@
To conclude, we want to point out that the constants in the bounds obtained from Lemma A.2, Lemma A.3 and Lemma A.4 can be chosen uniformly for all $T$ large enough and for $\gamma$ in a closed subset of $(0,1)$. In fact, as we have $\kappa=o(T^{1/3})$ and $\varkappa$ is contained in a bounded subset of $\mathbb{R}$, for large times the appearing densities $\rho_{\pm}=\sqrt{\gamma}+(\tfrac{1}{2}\varkappa\sqrt{\gamma}\pm\kappa)t^{-1/3}+{\cal O}(\varkappa^{2}t^{-2/3})$ are contained in a closed subset of $(0,1)$ as well. Therefore, the constants can be chosen uniformly for all those densities. Also the application of the exponential Chebyshev inequality for (3.37) allows the choice of uniform constants $C,c>0$. ∎
To see that the choice of $\kappa=o(T^{1/3})$ is admissible, refer to Remark A.5. In order to obtain uniform constants for different values of $\kappa$, we demand $\kappa={\cal O}(T^{\sigma})$ for some $\sigma\in(0,\tfrac{1}{3})$.
@@ 4.1 Control of fluctuations to the right
Since our notion of backwards paths shows not only similarities but also differences to the backwards geodesics in the cases of LPP models [4, 13] and height function representations [12], our implementation of the approach of [4] requires further technicalities, namely the use of Lemma 3.7. Moreover, it only gives us control of the right fluctuations of the backwards path. Also, due to the application of Proposition 3.16, we only get $1$ as exponent in the exponential decay, instead of $2$ respectively $3$ [13, 12]. The next proposition captures the control of right fluctuations of the backwards path.
@@
Proof.
We define
$l(t):=(1-2\sqrt{\gamma})t,\ m\coloneqq\min\{n\in\mathbb{N}|2^{-n}T\leq T^{1/2}\}$ (4.6)
and choose $u_{1}<u_{2}<\dots$ by $u_{1}\coloneqq\frac{K}{10}\textrm{ and }u_{n}-u_{n-1}=u_{1}2^{-(n-1)/2}$. We divide the interval $[0,T]$ up into $2^{n}$ subintervals for each $n\in\{1,\dots,m\}$ and denote
$A_{n}$$\coloneqq\{x_{N(T\downarrow k2^{-n}T)}(k2^{-n}T)\leq l(k2^{-n}T)+u_{n}T^{2/3},0\leq k\leq 2^{n}\},$ (4.7)
[…]
$L$$\coloneqq\Big{\{}\sup_{s\in[0,1]}|x_{N(T\downarrow(k+s)2^{-m}T)}((k+s)2^{-m}T)-x_{N(T\downarrow k2^{-m}T)}(k2^{-m}T)-l(s2^{-m}T)|$
$\qquad\leq\tfrac{K}{2}T^{2/3},0\leq k\leq 2^{m}-1\Big{\}}$
as well as
$G$$\coloneqq\Big{\{}\sup_{t\in[0,T]}\{x_{N(T\downarrow t)}(t)-l(t)\}\leq KT^{2/3}\Big{\}}.$ (4.8)
$F\coloneqq\{x_{N(T\downarrow k2^{-m}T)}(k2^{-m}T)>l(k2^{-m}T)+\tfrac{K}{2}T^{2/3}\textrm{ for some }k\in\{0,\dots,2^{m}-1\}\}.$ (4.9)
The first feature we need is the inclusion
Heuristically, (4.10) implies that if the backwards path fluctuates to the right at one edge point of the subintervals of length $2^{-m}T$, then this either happens at time 00 or at time $T$, or we find a decomposition of $[0,T]$ into larger dyadic subintervals such that at one of their inner edge points, the backwards path fluctuates to the right, while at the two adjacent edge points, it stays to the left of a suitable line.
$x_{N(T\downarrow k2^{-m}T)}(k2^{-m}T)>l(k2^{-m}T)+u_{m-n_{0}}T^{2/3}$ (4.11)
is fulfilled by some $k=2^{n_{0}}z$ with $z\in\{1,\dots,2^{m-n_{0}}-1\}$ odd. Then, $B_{m-n_{0},z}$ occurs. If $n_{0}=m-1$, then we have $B_{m-n_{0},z}=B_{1,1}$. Else, we show that $A_{m-n_{0}-1}^{c}$ taking place would contradict the maximality of $n_{0}$ and conclude that $B_{m-n_{0},z}\cap A_{m-n_{0}-1}$ occurs. Since we have $F^{c}\cap L\subseteq G$, the inclusion (4.10) implies
Bound on $\mathbb{P}(L^{c})$.
$|x_{N(T\downarrow(k+s)2^{-m}T)}((k+s)2^{-m}T)-x_{N(T\downarrow k2^{-m}T)}(k2^{-m}T)|$ (4.13)
is bounded by
$\mathcal{P}((k+s)2^{-m}T)-\mathcal{P}(k2^{-m}T).$ (4.14)
$\mathbb{P}(L^{c})\leq Ce^{-cKT^{2/3}}.$ (4.15)
Figure 4: The event $B_{n,k}\cap A_{n-1}$ for $k\in\{1,\dots,2^{n}-1\}$ odd.
Bound on $\sum_{n=2}^{m}\sum_{k=1}^{2^{n}-1}\mathbb{P}(B_{n,k}\cap A_{n-1})+\sum_{k=0}^{2}\mathbb{P}(B_{1,k})$.
$t_{1}\coloneqq(k-1)2^{-n}T,t_{2}\coloneqq(k+1)2^{-n}T,t\coloneqq k2^{-n}T$ (4.16)
and obtain by $A_{n-1}$:
$x_{N(T\downarrow t_{1})}(t_{1})$ $\leq l(t_{1})+u_{n-1}T^{2/3}\eqqcolon x_{1},$ (4.17)
$x_{N(T\downarrow t_{2})}(t_{2})$ $\leq l(t_{2})+u_{n-1}T^{2/3}\eqqcolon x_{2}.$ (4.18)
Suitable bounds on $\mathbb{P}(B_{1,0}),\mathbb{P}(B_{1,1})$ and $\mathbb{P}(B_{1,2})$ are obtained by Lemma A.2, see also Remark A.5, by Theorem 4.4 and lastly by Lemma A.1. For all bounds, the constants $C,c>0$ can be chosen uniformly for large $T$ and for $\gamma$ in a closed subset of $(0,1)$. Combining all bounds, we conclude $\mathbb{P}(G^{c})\leq Ce^{-cK}$. ∎
4.2 Mid-time estimate and its application
@@
Let $\tilde{N}\coloneqq\tfrac{\alpha}{2}T+2KT^{2/3}$. In the following, we show
@@
$\mathbb{P}(N(T\downarrow\tfrac{T}{2})\leq\tilde{N})\geq$ $\mathbb{P}\Bigl{(}x_{N}(T)<S<\min_{\tilde{N}<n\leq N}\{x_{n}(\tfrac{T}{2})+x^{\textrm{step}}_{N-n+1}(\tfrac{T}{2})\}\Bigr{)}$ (4.24)
@@
Choosing $S\coloneqq(1-2\sqrt{\alpha})T+\phi K^{2}T^{1/3}$ with $\phi\coloneqq\tfrac{1}{4\alpha^{3/2}}$, Lemma A.1 yields
@@
Further, we define $h(n)\coloneqq\tfrac{1}{\sqrt{\alpha}}(n-\tfrac{\alpha}{2}T)$ and observe
@@
Claim A: For all $T$ large enough, it holds (4.29)$\leq Ce^{-cK^{2}T^{2\delta}}$. The constants $C,c>0$ can be taken uniformly for $\alpha$ in a closed subset of $(0,1)$.
@@
($\leq$ $\sum_{KT^{2/3+\delta}<k\leq\tfrac{\alpha}{2}\xi_{0}T}\mathbb{P}(x_{{\alpha\over 2}T+k}(\tfrac{T}{2})\leq(1-2\sqrt{\alpha+2kT^{-1}})\tfrac{T}{2}-\tfrac{k^{2}}{4\alpha^{3/2}}T^{-1}+\tfrac{\phi}{2}K^{2}T^{1/3})$ (4.34)
@@
(4.28)$\leq\sum_{\ell=2}^{T^{\delta}-1}\mathbb{P}\Bigl{(}\min_{k\in I_{\ell}}\{x_{\tfrac{\alpha}{2}T+k}(\tfrac{T}{2})+h(\tfrac{\alpha}{2}T+k)\}\leq\tfrac{S}{2}\Bigr{)}$ ) (4.35)
for $I_{\ell}\coloneqq(\ell KT^{2/3},(\ell+1)KT^{2/3}]\cap\mathbb{Z}$. The $\ell$th term in (4.35) can be bounded by
@@
$+\mathbb{P}\Bigl{(}x_{\tfrac{\alpha}{2}T+\ell KT^{2/3}}(\tfrac{T}{2})+h(\tfrac{\alpha}{2}T+\ell KT^{2/3})\leq\tfrac{S}{2}+\tilde{\phi}\ell^{2}K^{2}T^{1/3}\Bigr{)},$ (4.37)
where we choose $\tilde{\phi}\coloneqq\tfrac{1}{3\alpha^{3/2}}$.
@@
We couple $x(t)$ and $x^{\rho_{-}}(t)$ by clock coupling and apply Proposition 3.16 for $t,N_{+}$ and $P_{-}$. Instead of $T$, we consider the time $\tfrac{T}{2}$. With this, $t=\tfrac{T}{2}$ leads to $\varkappa=0$, and we get $N_{+}=\gamma\tfrac{T}{2}$ for $\gamma=\alpha+2(\ell+1)KT^{-1/3}$. Since $(\ell+1)K\leq T^{\delta}K=o(T^{1/3})$ (as $\varepsilon+\delta<\tfrac{1}{3}$), $\gamma$ is contained in a closed subset of $(0,1)$ for $T$ large enough and we have $\kappa=o(T^{1/3})$. As seen in the proof of Proposition 3.16, it holds
@@
$q\coloneqq\rho-\sqrt{\alpha}=-\tfrac{KT^{-1/3}}{\sqrt{\alpha}}-\tfrac{(\ell+1)^{2}K^{2}T^{-2/3}}{2\alpha^{3/2}}+{\cal O}(\ell^{3}K^{3}T^{-1})$ (4.46)
@@
Conclusion.
@@
Now, we return to the setting in the proof of Proposition 4.2. Recall that we want to bound
@@
[…]
@@
The explicit choice of the parameter $\phi_{n}$ becomes clear in the proof of Proposition 4.6.
@@
$\mathbb{P}(M>\tilde{M})$ $=\mathbb{P}(x^{\textrm{step},x_{1}}_{\tilde{M}}(t_{1},t_{2})>x_{2})$ (4.60)
$=\mathbb{P}(x^{\textrm{step}}_{\tilde{M}}(t_{2}-t_{1})>x_{2}-x_{1})$
@@
$\leq\mathbb{P}(x^{\textrm{step},x_{1}}_{\tilde{M}(t_{2}\downarrow t)}(t_{1},t)>l(t)+u_{n}T^{2/3})+\mathbb{P}(M\leq\tilde{M})$
$\leq\mathbb{P}(x^{\textrm{step}}_{\tilde{M}(t_{2}-t_{1}\downarrow{{t_{2}-t_{1}}\over 2})}(\tfrac{t_{2}-t_{1}}{2})>l(t)-x_{1}+u_{n}T^{2/3})+\mathbb{P}(M\leq\tilde{M}),$
where the second inequality holds by Lemma 3.7 and for the forth inequality, we use
@@
since $t=\tfrac{t_{1}+t_{2}}{2}$. Furthermore, by Lemma 4.5, (4.63) is bounded from above by
@@ 4.3 Control of fluctuations to the left
${\cal C}^{r}$ =$=\{(x,t)\in\mathbb{Z}\times[0,T]\ |\ x\leq(1-2\sqrt{\gamma})t+KT^{2/3}\},$ (4.71)
${\cal C}^{l}$ =$=\{(x,t)\in\mathbb{Z}\times[0,T]\ |\ x\geq(1-2\sqrt{\gamma})t-KT^{2/3}\},$
@@
Then, Lemma 3.1 implies
@@
It remains to bound $\mathbb{P}(E_{M,{\cal C}^{l}})$ from below. We choose the label $M\in\mathbb{N}$ such that the event $\{x^{h}_{M}(T)<x_{N}(T)\}$ has a high probability, but the macroscopic position of the backwards path $x^{h}_{M(T\downarrow t)}(t)$ is still within an $o(T^{2/3})-$neighbourhood of the macroscopic position of the backwards path $x_{N(T\downarrow t)}(t)$. Since $N=\gamma T$ with $\gamma\in(0,1)$, we choose $M=(1-\sqrt{\gamma})^{2}T-T^{1/2}$, see Figure 6.
@@
Then, Lemma A.1 yields

### v2

@@ 3.2 Asymptotic results
$\mathbb{P}(N(t\downarrow 0)>\kappa t^{1/3})=\mathbb{P}(|x_{N(t\downarrow 0)}(0)|\geq\kappa t^{1/3})\leq Ce^{-c\kappa}$ (3.34)
and
$\mathbb{P}(x_{P}^{\rho_{-}}(t)\leq(1-2\rho_{-})t-\kappa t^{2/3})\geq 1-Ce^{-c\kappa},$ (3.35)
@@
$Z_{\kappa,t}:=\#\{\textrm{particles in }x^{\rho_{-}}(0)\textrm{ at sites in }\{-\tfrac{1}{2}\kappa t^{2/3},\dots,-1\}\}\sim\textrm{Bin}(\tfrac{1}{2}\kappa t^{2/3},\rho_{-}).$ (3.37)
@@
with the probability greater than $1-Ce^{-c\kappa}$.
@@
To conclude, we want to point out that the constants in the bounds obtained from Lemma A.2, Lemma A.3 and Lemma A.4 can be chosen uniformly for all $T$ large enough and for $\gamma$ in a closed subset of $(0,1)$. In fact, as we have $\kappa=o(T^{1/3})$ and $\varkappa$ is contained in a bounded subset of $\mathbb{R}$, for large times the appearing densities $\rho_{\pm}=\sqrt{\gamma}+(\tfrac{1}{2}\varkappa\sqrt{\gamma}\pm\kappa)t^{-1/3}+{\cal O}(\varkappa^{2}t^{-2/3})$ are contained in a closed subset of $(0,1)$ as well. Therefore, the constants can be chosen uniformly for all those densities. Also the application of the exponential Chebyshev inequality for (3.38) allows the choice of uniform constants. ∎
To see that the choice of $\kappa=o(T^{1/3})$ is admissible, refer to Remark A.5.
@@ 4.1 Control of fluctuations to the right
Our notion of backwards paths shows not only similarities but also differences to the backwards geodesics in the cases of LPP models [4, 13] and height function representations [12]. In particular, our implementation of the approach of [4] requires the use of Lemma 3.7. This, however, only gives us control of the right fluctuations of the backwards path. Also, due to the application of Proposition 3.16, we only get $1$ as exponent in the exponential decay, instead of $2$ respectively $3$ [13, 12]. The next proposition captures the control of right fluctuations of the backwards path.
@@
4.2 Mid-time estimate
@@
Let $\tilde{N}:=\tfrac{\alpha}{2}T+2KT^{2/3}$. In the following, we show
@@
[…]
@@
Choosing $S:=(1-2\sqrt{\alpha})T+\phi K^{2}T^{1/3}$ with $\phi:=\tfrac{1}{4\alpha^{3/2}}$, Lemma A.1 yields
@@
Further, we define $h(n):=\tfrac{1}{\sqrt{\alpha}}(n-\tfrac{\alpha}{2}T)$ and observe
@@
[…]
@@
[…]
@@
[…]
for $I_{\ell}:=(\ell KT^{2/3},(\ell+1)KT^{2/3}]\cap\mathbb{Z}$. The $\ell$-th term in (4.20) can be bounded by the sum of
@@
and
$\mathbb{P}\Bigl{(}x_{\tfrac{\alpha}{2}T+\ell KT^{2/3}}(\tfrac{T}{2})+h(\tfrac{\alpha}{2}T+\ell KT^{2/3})\leq\tfrac{S}{2}+\tilde{\phi}\ell^{2}K^{2}T^{1/3}\Bigr{)},$ (4.22)
where we choose $\tilde{\phi}:=\tfrac{1}{3\alpha^{3/2}}$.
@@
with uniform constants for all $T$ large enough, $\alpha$ in a closed subset of $(0,1)$ and $\ell\in\{2,\dots,T^{\delta}-1\}$. Summing over $\ell$, Claim B.1 follows.
@@
We couple $x(t)$ and $x^{\rho_{-}}(t)$ by clock coupling and apply Proposition 3.16 for $t,N_{+}$ and $P_{-}$. Instead of $T$, we consider the time $\tfrac{T}{2}$. With this, $t=\tfrac{T}{2}$ leads to $\varkappa=0$, and we get $N_{+}=\gamma\tfrac{T}{2}$ for $\gamma=\alpha+2(\ell+1)KT^{-1/3}$. Since $(\ell+1)K\leq T^{\delta}K=o(T^{1/3})$ (as $\varepsilon+\delta<\tfrac{1}{3}$), $\gamma$ is contained in a closed subset of $(0,1)$ for $T$ large enough and we have $\kappa=o(T^{1/3})$. As seen in the proof of Proposition 3.16 (see (3.40) and the paragraph below), it holds
@@
$q:=\rho-\sqrt{\alpha}=-\tfrac{KT^{-1/3}}{\sqrt{\alpha}}-\tfrac{(\ell+1)^{2}K^{2}T^{-2/3}}{2\alpha^{3/2}}+{\cal O}(\ell^{3}K^{3}T^{-1})$ (4.31)
@@
With the bounds in Claim B.1 and Claim B.2, the proof of Claim B is completed.
Conclusion
@@
With the mid-time estimate at hand, we now turn to the proof of Proposition 4.2.
4.3 Proof of Proposition 4.2
We define
$l(t):=(1-2\sqrt{\gamma})t,\ m:=\min\{n\in\mathbb{N}|2^{-n}T\leq T^{1/2}\}$ (4.41)
and choose $u_{1}<u_{2}<\dots$ by $u_{1}:=\frac{K}{10}\textrm{ and }u_{n}-u_{n-1}=u_{1}2^{-(n-1)/2}$. We divide the interval $[0,T]$ up into $2^{n}$ subintervals for each $n\in\{1,\dots,m\}$ and denote
$A_{n}$ $=\{x_{N(T\downarrow k2^{-n}T)}(k2^{-n}T)\leq l(k2^{-n}T)+u_{n}T^{2/3},0\leq k\leq 2^{n}\},$ (4.42)
$B_{n,k}$ $=\{x_{N(T\downarrow k2^{-n}T)}(k2^{-n}T)>l(k2^{-n}T)+u_{n}T^{2/3}\},0\leq k\leq 2^{n},$
as well as
$G$ $=\Big{\{}\sup_{t\in[0,T]}\{x_{N(T\downarrow t)}(t)-l(t)\}\leq KT^{2/3}\Big{\}}.$ (4.43)
$F:=\{x_{N(T\downarrow k2^{-m}T)}(k2^{-m}T)>l(k2^{-m}T)+\tfrac{K}{2}T^{2/3}\textrm{ for some }k\in\{0,\dots,2^{m}-1\}\}.$ (4.44)
The first feature we need is the inclusion
Heuristically, (4.45) implies that if the backwards path fluctuates to the right at one edge point of the subintervals of length $2^{-m}T$, then this either happens at time 0 or at time $T$, or we find a decomposition of $[0,T]$ into larger dyadic subintervals such that at one of their inner edge points, the backwards path fluctuates to the right, while at the two adjacent edge points, it stays to the left of a suitable line.
$x_{N(T\downarrow k2^{-m}T)}(k2^{-m}T)>l(k2^{-m}T)+u_{m-n_{0}}T^{2/3}$ (4.46)
is fulfilled by some $k=2^{n_{0}}z$ with $z\in\{1,\dots,2^{m-n_{0}}-1\}$ odd. Then, $B_{m-n_{0},z}$ occurs. If $n_{0}=m-1$, then we have $B_{m-n_{0},z}=B_{1,1}$. Else, we show that $A_{m-n_{0}-1}^{c}$ taking place would contradict the maximality of $n_{0}$ and conclude that $B_{m-n_{0},z}\cap A_{m-n_{0}-1}$ occurs. This shows (4.45).
Denoting
$L$ $=\Big{\{}\sup_{s\in[0,1]}|x_{N(T\downarrow(k+s)2^{-m}T)}((k+s)2^{-m}T)-x_{N(T\downarrow k2^{-m}T)}(k2^{-m}T)-l(s2^{-m}T)|$ (4.47)
$\qquad\leq\tfrac{K}{2}T^{2/3},0\leq k\leq 2^{m}-1\Big{\}},$
we have $F^{c}\cap L\subseteq G$ and the inclusion (4.45) implies
Bound on $\mathbb{P}(L^{c})$
$\sup_{s\in[0,1]}|x_{N(T\downarrow(k+s)2^{-m}T)}((k+s)2^{-m}T)-x_{N(T\downarrow k2^{-m}T)}(k2^{-m}T)|$ (4.49)
is bounded by
$\mathcal{P}((k+1)2^{-m}T)-\mathcal{P}(k2^{-m}T).$ (4.50)
$\mathbb{P}(L^{c})\leq Ce^{-cKT^{2/3}}.$ (4.51)
Figure 4: The event $B_{n,k}\cap A_{n-1}$ for $k\in\{1,\dots,2^{n}-1\}$ odd.
Bound on $\sum_{n=2}^{m}\sum_{k=1}^{2^{n}-1}\mathbb{P}(B_{n,k}\cap A_{n-1})+\sum_{k=0}^{2}\mathbb{P}(B_{1,k})$
$t_{1}:=(k-1)2^{-n}T,t_{2}:=(k+1)2^{-n}T,t:=k2^{-n}T$ (4.52)
and obtain by $A_{n-1}$:
$x_{N(T\downarrow t_{1})}(t_{1})$ $\leq l(t_{1})+u_{n-1}T^{2/3}=:x_{1},$ (4.53)
$x_{N(T\downarrow t_{2})}(t_{2})$ $\leq l(t_{2})+u_{n-1}T^{2/3}=:x_{2}.$ (4.54)
Suitable bounds on $\mathbb{P}(B_{1,0}),\mathbb{P}(B_{1,1})$ and $\mathbb{P}(B_{1,2})$ are obtained by Lemma A.2, see also Remark A.5, by Theorem 4.4 and lastly by Lemma A.1.
For all bounds, the constants $C,c>0$ can be chosen uniformly for large $T$ and for $\gamma$ in a closed subset of $(0,1)$. Combining all bounds, we conclude $\mathbb{P}(G^{c})\leq Ce^{-cK}$.
4.4 Application of a comparison inequality and the mid-time estimate
It remains to bound
@@ 4.4 Application of a comparison inequality and the mid-time estimate
Define $\phi_{n}:=\sqrt{\gamma}\tfrac{1}{20}(2^{n-1})^{1/6}\ \textrm{and}\ \tilde{M}:=\gamma 2^{-n+1}T-\phi_{n}K(2^{-n+1}T)^{2/3}.$ Then, it holds
@@
[…]
@@
[…]
where the second inequality holds by Lemma 3.7 and for the fourth inequality, we use
@@
since $t=\tfrac{t_{1}+t_{2}}{2}$. Furthermore, by Lemma 4.5, the right hand side of (4.64) is bounded from above by
@@ 4.5 Control of fluctuations to the left
${\cal C}^{l}$ $=\{(x,t)\in\mathbb{Z}\times[0,T]\ |\ x\geq(1-2\sqrt{\gamma})t-KT^{2/3}\},$
@@
Recall that the process $\tilde{x}^{r}(t)$ with $\tilde{x}^{r}(0)=x(0)$ is constructed by the same Poisson events as $x(t)$ in ${\cal C}^{r}$ and with $({\cal C}^{r})^{c}$ completely filled by holes. Likewise, $\tilde{x}^{l}(t)$ is the process with $\tilde{x}^{l}(0)=x(0)$, constructed by the same Poisson events as $x(t)$ in ${\cal C}^{l}$ and with $({\cal C}^{l})^{c}$ completely filled by particles. Then, Lemma 3.1 implies
@@
It remains to bound $\mathbb{P}(E_{M,{\cal C}^{l}})$ from below for some suitable $M$. We choose the label $M\in\mathbb{N}$ such that the event $\{x^{h}_{M}(T)<x_{N}(T)\}$ has a high probability, but the macroscopic position of the backwards path $x^{h}_{M(T\downarrow t)}(t)$ is still within an $o(T^{2/3})$-neighbourhood of the macroscopic position of the backwards path $x_{N(T\downarrow t)}(t)$. Since $N=\gamma T$ with $\gamma\in(0,1)$, we choose $M=(1-\sqrt{\gamma})^{2}T-T^{1/2}$, see Figure 6.
@@
Then, Lemma A.1 (with $\alpha=\gamma$ and $c_{1}(\alpha)s=T^{1/6}$) yields